Sodium metal crystalizes in the BCC structure with an edge length of unit cell 

equal to 4.29 x 10 -8

cm. Calculate the atomic radius of sodium metal.


1
Expert's answer
2021-10-29T02:30:11-0400

Since, in the question it is given that the edge length a of the unit cell is given by 4.29*10^-8 which can be written in 4.29 angstroms. Hence, we get that the value of r after putting all the values as = √3 *4.29/4 = 1.73 * 4.29 angstroms/4 = 1.86 angstroms.